# PET-MAD, a universal interatomic potential for advanced materials modeling

### v1

@@ II.3 Benchmarking
The benchmarking results are presented in Table 1. We find that PET-MAD achieves high accuracy in predicting both energies and interatomic forces, outperforming MACE-MP-0-L in most cases, and competing closely with MatterSim-5M and SevenNet-l3i5 even for the inorganic datasets they are trained on. Orb-v2 performs significantly better than all other models on MPtrj and Alexandria, which are part of tis training set. PET-MAD outperforms all other models for molecular data sets such as SPICE and MD22, and matches Orb-v2 also on the OC2020 S2EF – which is a truly extrapolative exercise as MAD does not include configurations of adsorbed molecules. As discussed above, MACE-MP-0-L, MatterSim-5M, Orb-v2 and SevenNet-l3i5 perform significantly worse than PET-MAD on the MAD benchmark. Importantly the PET-MAD model achieves state-of-the-art performance while being computationally efficient. We compared the throughput of molecular dynamics simulations of systems of different sizes and densities (solid aluminum, diamond, and liquid water, see Figure 2) and, when available, used lighter versions of the reference models compared to those used in the accuracy benchmarks (namely, MACE-MP-0 M, MatterSim-1M, and SevenNet-0). These models have fewer parameters and thus allow faster inference at the expense of accuracy, making this a more challenging test for the speed of PET-MAD. Nevertheless, we see that PET-MAD is faster and more memory efficient than all models except Orb-v2, which benefits from the faster inference provided by non-conservative force prediction.
Overall, our benchmarking results highlight the accuracy and speed of PET-MAD across several classes of materials. Its ability to generalize well across both organic and inorganic domains, energy conservation and well-learned rotational equivariance makes it preferable for advanced atomistic simulations. In the following sections, we put it to even more stringent tests, demonstrating its applicability in various simulation scenarios. For each example, we re-computed with DFT setting corresponding to those used for the MAD dataset the structures used in previous studies, and train a bespoke PET model that we use as the ground truth, and a benchmark for the generally-applicable PET-MAD and a version fine-tuned to each ad hoc dataset.
@@ II.4 Ionic transport in lithium thiophosphate
We compare the results obtained using the PET-MAD potential with a single-purpose PET potential trained from scratch over the dataset from Ref. 42 (PET-Bespoke), and with one fine-tuned over the same dataset using LoRA technique (see Section IV.5 for more details, and the SI for a full discussion). For this dataset, the validation mean absolute error (MAE) of PET-MAD for energy| forces is 4.9 meV/atom|63.9 meV/Å, to be compared with 1.2meV/atom|35.6 meV/Å, for a model trained from scratch on the LPS dataset, and 1.4 meV/atom|39.2 meV/Å  for the fine-tuned model. Figure 3 shows the values of the conductivity for all the phases and over a wide range of temperatures computed via PET-MAD (orange), a model fine-tuned using the LoRA algorithm with rank 8 (light blue) and the PET-Bespoke model (blue). For the $\alpha$ and $\beta$ phases, PET-MAD is in excellent agreement with the bespoke and fine-tuned models, demonstrating quantitative accuracy despite the larger validation error. For the $\gamma$-phase, PET-MAD slightly overestimates the temperature at which the transition to a high-conductivity phase occurs, as well as the values of $\sigma$ for $T<625$. Except for these small errors, PET-MAD has excellent agreement with the more accurate dedicated models, capturing the change of behavior in $\sigma$ following the rotation of the PS4 tetrahedra [42]. The results are also in quantitative agreement with those of Ref. 42, even though they were based on slightly different DFT parameters and relied on a different class of MLIPs.
Figure 4: Chemical potential differences between liquid and solid GaAs phases as a function of temperature computed with PET-MAD (orange), PET-Bespoke (blue) and PET-MAD-LoRA (light blue) models. Dashed solid lines indicate mean potential differences and light colored lines individual committee member reweighted potential differences - the large spread of the PET-MAD ensemble predictions correctly captures the prediction uncertainty with respect to the Bespoke and LoRA finetuned model.
@@ II.5 Melting point of GaAs
Gallium Arsenide (GaAs) is a III-V semiconductor whose properties make it a promising candidate for new optical and electrical devices, as well as for high-end photovoltaics. The growth of GaAs nanostructures often relies on the coexistence of solid GaAs and molten Ga in an As-rich atmosphere,[46], requiring MLIPs which can describe accurately different phases and compositions across the Ga/As phase diagram. Electronic properties change wildly between different phases, posing additional challenges to the construction of empirical potentials. Following the work of Imbalzano et al. [47], we repeat the calculation of the melting point of stoichiometric GaAs at ambient pressure with the interface pinning method[48] effectively estimating the phase diagram of GaAs along an isopleth at fixed pressure. The method computes the chemical potential differences between the liquid and solid phases, and the melting point is determined identifying the temperature at which the chemical potential difference becomes zero. This also allows us to demonstrate a feature of our PET implementation, that uses a Laplace approximation[28] to generate an ensemble of potentials for uncertainty quantification. In combination with reweighing techniques, this allows us to propagate errors on the chemical potential curves[49], and hence assess the contribution to the error on the melting point that comes from the limitations of the ML fit (the epistemic part of the error).
@@
Figure 4 demonstrates these calculations, comparing the PET-MAD model with a bespoke PET model trained on the reference GaAs training set (Bespoke) and the LoRA fine-tuned model on the same training set. The test set errors (MAE) on energies| forces are 14.4 meV/at.|74.1 meV/Å  for PET-MAD, 0.7 meV/at.|29.0 meV/Å  for PET-Bespoke and 1.3 meV/at.|45.3 meV/Å  for PET-MAD-LoRA. When it comes to the estimation of the melting point, the LoRA and Bespoke models are in quantitative agreement ($1169\pm 3$ K vs $1169\pm 4$ K) whilst the PET-MAD model predicts a slightly smaller value of $1111\pm 72$ K. The predicted error is consistent with the lower accuracy of the general-purpose PET-MAD, which highlights the importance of the inexpensive uncertainty propagation afforded by the Laplace ensemble architecture implemented in our PET models. It should be noted that the error of the general-purpose model against the bespoke potentials is very small in comparison with the deviation of the computed melting point from the experimentally measured melting point (1511 K), and with the typical errors of empirical forcefields. This large discrepancy is consistent with that observed in Ref. 47, and can be attributed to shortcomings of the reference electronic-structure calculations – which is not uncommon as the computed melting point of a given material can vary by hundreds of Kelvin depending on the choice of DFT functional[50].
@@
Figure 5: Gibbs surface excess $\Gamma_{a}$ of the elements on a (111) surface of the CoCrFeMnNi alloy at 800 K from the REMD/MC simulation performed with the pre-trained PET-MAD model (orange bars), bespoke PET model trained on a subset of the HEA25S dataset (dark-blue bars), and the PET-MAD-LoRA model, fine-tuned on the same subset (light-blue bars).
@@ II.6 Surface segregation in high-entropy alloys
High-entropy alloys – containing 5 or more metals in roughly equimolar composition – often possess excellent mechanical properties[51, 52], and perform well as heterogeneous catalysts[53, 54]. Investigating their properties, and exploring efficiently their composition space, requires versatile models that can handle a high degree of chemical diversity. We replicate some of the simulations in Ref. 55, using PET-MAD to study the segregation of different elements at the (111) surface of the CoCrFeMnNi alloy, a prototypical HEA[52]. Modeling of the differential surface propensity of the elements in the alloy – which is central to the applications of HEAs as heterogeneous catalysts – requires a combination of replica-exchange molecular dynamics with Monte-Carlo atom swaps (REMD/MC) to overcome the slow diffusivity, and the existence of free-energy barriers to segregation starting from a random alloy. This example allows us to demonstrate the capabilities of PET-MAD in complex computational workflows, to assess the ability of the model to describe surface effects that are often overlooked in other universal models [56], and to test fine-tuning over a dataset with a greater degree of chemical diversity than the other examples we consider. We use the Gibbs surface excess calculated at 800 K using the REMD/MC protocol (see Section S8.3 for computational details) as a measure of surface segregation in a CoCrFeMnNi surface slab, We compare three models: the pre-trained PET-MAD, PET trained from scratch (PET-Bespoke) and the fine-tuned model (PET-MAD-LoRA), where the last two were trained on a subset of the HEA25S dataset [55], recomputed with MAD dataset DFT settings for consistency (details on the subset selection are given in the Supplementary Section S8.3). The validations MAEs of all three models on energy|forces are 25.8 meV/atom|175.1 meV/Å, 14.6 meV/atom|138.3 meV/Å, and 9.4 meV/atom|124.8 meV/Å, respectively. We compare three models: the pre-trained PET-MAD, PET trained from scratch (PET-Bespoke) and the fine-tuned model (PET-MAD-LoRA), where the last two were trained on a subset of the HEA25S dataset [55], recomputed with MAD dataset DFT settings for consistency (details on the subset selection are given in the Supplementary Section S8.3). The validations MAEs of all three models on energy| forces are 25.8 meV/atom|175.1 meV/Å, 14.6 meV/atom|138.3 meV/Å, and 9.4 meV/atom|124.8 meV/Å, respectively.
@@
Results on surface excess calculation are given in Figure 5. First, we note that both pre-trained and fine-tuned models yield almost exactly the same segregation pattern, which matches quantitatively with the results from Ref. 55, corresponding to surface layers enriched in nickel and depleted in other elements. Both these models agree qualitatively with the results of the PET-Bespoke model, but there is a more pronounced quantitative difference in the surface excess values. In interpreting this discrepancy, one should keep in mind that, contrary to other examples, we only recomputed about 2000 structures out of the 30’000 in the HEA25S, and so a bespoke model incurs a high risk of overfitting – which is also consistent with the learning curves reported in the SI. This case study demonstrates that pre-trained PET-MAD is capable of capturing all effects relevant to surface segregation in HEAs and providing advanced sampling capabilities of the HEA chemical and structural space without being specifically trained on HEA data.
@@
Figure 6: O-O, O-H radial distribution functions $g$ and heat capacity $C_{V}$ of liquid water at 298 K and 1 atm, calculated using PET-MAD, a fine-tuned version of PET-MAD, and a bespoke PET model, as a function of the number of beads employed in the PIMD simulation. q-TIP4P/F results are also included (for the heat capacity, these are reproduced from Ref. 57).
@@ II.7 Quantum nuclear effects in liquid water
Water is widely studied because of its biological, environmental and technological significance, as well as the many anomalies in its physical properties, which can be traced to the strong hydrogen bonds that form local tetrahedral motifs. Due to its light hydrogen atoms, accounting for nuclear quantum effects (NQEs) is fundamental in order to extract accurate observables for liquid water, even at room temperature[58]. In this example, we employ path integral molecular dynamics (PIMD) to model nuclear quantum effects in the simulation of a medium-sized periodic system consisting of 128 water molecules (384 atoms).
We evaluate the O-O pair correlation function, reporting on inter-molecular structural correlations the O-H pair correlation function, showing the extent of intra-molecular quantum fluctuations, and the constant-volume heat capacity at 298 K, a physical quantity that exhibits very pronounced quantum effects, and that requires the use of sophisticated path integral estimators that are notoriously difficult to converge[59]. We compare (1) the PET-MAD model, (2) a system-specific PET model trained on the dataset from Ref. 60, with energies and forces recomputed at the same level of theory used for MAD, (3) a PET-MAD model which was fine-tuned on the same water dataset and (4) values for a simple classical forcefield that reproduces well most of the experimental properties of water and allows us to demonstrate the slow convergence of some of the estimators [61].

### v2

@@ II.3 Benchmarking
The benchmarking results are presented in Table 1. We find that PET-MAD achieves high accuracy in predicting both energies and interatomic forces, outperforming MACE-MP-0-L in most cases, and competing closely with MatterSim-5M and SevenNet-l3i5 even for the inorganic datasets they are trained on. Upon using a consistent level of DFT theory, PET-MAD outperforms all other models in predicting raw energies of the WBM crystals from the Matbench Discovery benchmark [34]. Orb-v2 performs significantly better than all other models on MPtrj and Alexandria, which are part of its training set. PET-MAD outperforms all other models on molecular datasets such as SPICE and MD22, and matches Orb-v2 also on the OC2020 S2EF – which is a strongly extrapolative exercise as MAD does not include configurations of adsorbed molecules.
Figure 2: Comparison of accuracies of PET-MAD (black line), MACE-MP-0-L (blue line), Orb-v2 (orange line), MatterSim-5M (green line), and SevenNet-l3i5 (red line) models on various subsets of the MAD dataset. The left and center panels show the mean absolute errors in energies and force predictions. The right panel shows the rotational discrepancy in energy predictions of the unconstrained models (PET-MAD and Orb-v2). Results for all models except PET-MAD were obtained using consistently recalculated reference values (i.e. computed with MPtrj dataset settings) to achieve consistency with the training sets of the corresponding reference models. See also Table 1 for benchmarks on out-of-sample datasets.
The reference models perform much worse than PET-MAD on the (consistently recomputed) MAD benchmark. This is not too surprising, as MAD is designed to be more diverse than the datasets, on which these models are trained. Still, it is instructive to assess separately the accuracy of the various models on the different sections of MAD (Figure 2). PET-MAD outperforms MACE-MP-0-L, MatterSim-5M and SevenNet-l3i5 in almost all cases. The accuracy is similar (and in a few cases slightly better) for the stable inorganic structures including bulk (MC3D) and layered (MC2D) compounds. MC3D-derived surfaces and clusters show slightly larger force and considerably larger energy errors. Errors of the other universal models are up to 50 times larger for the distorted (MC3D-rattled and MC3D-randcomp) that contain especially unusual, diverse configurations. MatterSim-5M, which is trained on a broader, yet unpublished, set of configurations, performs substantially better than the other models. The most notable difference occurs in the molecular systems (SHIFTML-molcrys and SHIFTML-molfrags), where PET-MAD dramatically outperforms all other models. This result is again expected, as organic systems represent a completely different region of configuration space, which is heavily undersampled in the case of inorganic crystal datasets. The last model from the reference list, Orb-v2, requires a separate discussion, as it provides by default direct, non-conservative force predictions. Orb-v2 outperforms all other models (including, marginally, PET-MAD) on relaxed inorganic systems, presumably thanks to the combination of a flexible, unconstrained architecture and the large training set (which includes both MPtrj [15] and Alexandria [20] datasets). However, it still provides worse accuracy on rattled, random composition structures, surfaces and molecular systems compared to PET-MAD. Both PET-MAD and Orb-v2 use architectures that do not enforce exact rotational equivariance, which is learned approximately by data augmentation. The resulting symmetry breaking can be monitored and controlled easily [38], but it is important to assess its extent. To do so, we estimated the symmetry error in PET-MAD and Orb-v2 energy predictions by applying a series of rotations on a Lebedev-Laikov grid of order 9 for each structure in the MAD subset and calculating the standard deviation in the predictions (Figure 2, right panel). In most cases, the rotational discrepancy of the PET-MAD model is one or two orders of magnitude smaller than the corresponding prediction error, below 1 meV/atom for all subsets except MC3D-random and MC3D-cluster. Orb-v2 shows significantly higher discrepancies in energy predictions, which are at times comparable to the actual errors.
Figure 3: Inference time of several models evaluated over different bulk materials and system sizes. For each MLIP, we use its LAMMPS [39] interface if available, preferably choosing the Kokkos-enabled [40] ($k$) version, or its ASE [41] interface otherwise. All timings were measured on a single NVIDIA H100 GPU. Missing points for MACE-MP-0 and SevenNet caused by out-of-memory errors are marked with red crosses. The non-conservative (NC) versions of Orb-v2 and PET-MAD are shown in dashed lines. These models benefit from a theoretical speedup, but can violate conservation of energy, often resulting in ill-behaved molecular dynamics [33]. The model versions used were MatterSim-v1.0.0-1M, MACE-MP-0 (M), Orb-v2, SevenNet-0 (11Jul2024).
The PET-MAD model is also computationally efficient, which facilitates simulating larger time and length scales. We compared the throughput of molecular dynamics simulations of systems of different sizes and densities (solid aluminum, diamond, and liquid water, see Figure 3) and, when available, used lighter versions of the reference models compared to those used in the accuracy benchmarks (namely, MACE-MP-0 M, MatterSim-1M, and SevenNet-0). These models have fewer parameters and thus allow faster inference at the expense of accuracy, making this a more challenging test for the speed of PET-MAD. Nevertheless, we see that PET-MAD is faster and more memory efficient than all models. When exploiting the computational advantage of direct-force prediction, PET-MAD is also competitive with the non-conservative Orb-v2 model.
The takeaway is that an unconstrained architecture and a problem-agnostic construction of the training set make PET-MAD competitive in speed and accuracy with models that have a larger number of parameters and are trained on much larger datasets (Fig. 1). This has important implications to guide the design of future datasets, and provides a more sustainable and accessible platform to support model development.
Benchmarking accuracies, however, do not solely ensure that PET-MAD can be reliably used in realistic atomistic modeling scenarios. Therefore, in the following sections, we demonstrate applicability of PET-MAD for atomistic simulations by comparing its performance against single-purpose and finetuned models for six diverse and challenging use cases.
@@ II.4 Ionic transport in lithium thiophosphate
We compare the results obtained using the PET-MAD potential with a single-purpose PET potential trained from scratch over the dataset from Ref. 43 (PET-Bespoke), and with one finetuned over the same dataset using LoRA technique (see Section IV.5 for more details, and the SI for a full discussion). For this dataset, the validation mean absolute error (MAE) of PET-MAD for energy| forces is 4.9 meV/atom|63.9 meV/Å, to be compared with 1.2 meV/atom|35.6 meV/Å, for a model trained from scratch on the LPS dataset, and 1.3 meV/atom|36.0 meV/Å  for the finetuned model. Figure 4 shows the values of the conductivity for all the phases and over a wide range of temperatures computed with PET-MAD (black), a model finetuned using the LoRA algorithm with rank 8 (orange) and the PET-Bespoke model (blue). For all three phases, PET-MAD is in excellent agreement with the bespoke and finetuned models, demonstrating quantitative accuracy despite the larger validation error. For the $\gamma$-phase, PET-MAD slightly overestimates the temperature at which the transition to a high-conductivity phase occurs. Except for this minor discrepancy, PET-MAD shows excellent agreement with the more accurate dedicated models, capturing the change of behavior in $\sigma$ following the rotation of the PS4 tetrahedra [43]. The results agree quantitatively with those of Ref. 43, despite being based on slightly different DFT parameters and relied on a different class of MLIPs.
Figure 5: Chemical potential differences between liquid and solid GaAs phases as a function of temperature computed with PET-MAD (black), PET-Bespoke (blue) and PET-MAD-LoRA (orange) models. The dashed solid lines indicate the mean values of the predicted $\Delta\mu$, and the light colored lines represent the reweighted predictions from the individual members of the shallow ensemble. The large spread of the PET-MAD ensemble predictions correctly reflects the discrepancy with respect to the Bespoke and LoRA-finetuned model.
@@ II.5 Melting point of GaAs
Gallium Arsenide (GaAs) is a III-V semiconductor whose properties make it a good choice to manufacture the high-quality optical and electrical devices, as well as the high-end photovoltaics. The growth of GaAs nanostructures often relies on the coexistence of solid GaAs and molten Ga in an As-rich atmosphere [46]. Corresponding computational studies typically require MLIPs that can accurately describe different phases and compositions across the Ga/As phase diagram. Electronic properties change wildly between different phases, posing additional challenges to the construction of empirical potentials. Following the work of Imbalzano et al. [47], we repeat the calculation of the melting point of stoichiometric GaAs at ambient pressure with the interface pinning method. This method computes the differences in chemical potentials between the liquid and solid phases, and the melting point is determined by identifying the temperature at which the chemical potential difference becomes zero. A combination of the Laplace ensemble predictions with the reweighting technique allows us to propagate errors on the chemical potential curves[48], and therefore assess the contribution of the epistemic part of the error to the melting point error caused by limitations of the ML fit.
@@
Figure 5 demonstrates these calculations, comparing the PET-MAD model with a bespoke PET model trained on the reference GaAs training set, and the LoRA-finetuned model on the same training set. The test set errors (MAE) on energies| forces are 14.4 meV/at.|74.1 meV/Å  for PET-MAD, 0.7 meV/at.|29.0 meV/Å  for PET-Bespoke and 1.3 meV/at.|45.3 meV/Å  for PET-MAD-LoRA. When it comes to the estimation of the melting point, the LoRA and Bespoke models are in quantitative agreement ($1169\pm 3$ K vs $1169\pm 4$ K) whilst the PET-MAD model predicts a slightly smaller value of $1111\pm 72$ K. The predicted error is consistent with the lower accuracy of the general-purpose PET-MAD, which highlights the importance of the inexpensive uncertainty propagation afforded by the Laplace ensemble architecture implemented in our PET models. It should be noted that the error of the general-purpose model against the bespoke potentials is very small compared to the deviation of the computed melting point from the experimentally measured melting point (1511 K), and with the typical errors of empirical forcefields. This large discrepancy is consistent with that observed in Ref. 47, and can be attributed to shortcomings of the reference electronic-structure calculations – which is not uncommon as the computed melting point of a given material can vary by hundreds of Kelvin depending on the choice of DFT functional[49].
@@
Figure 6: Gibbs surface excess $\Gamma_{a}$ of the elements on a (111) surface of the CoCrFeMnNi alloy at 800 K from the REMD/MC simulation performed with the pre-trained PET-MAD model (black bars), bespoke PET model trained on a subset of the HEA25S dataset (blue bars), and the PET-MAD-LoRA model, finetuned on the same subset (orange bars). Both pre-trained and finetuned models result in almost identical segregation pattern with the surface enriched in nickel, which is similar to one from Ref. 50. The inconsistency between the bespoke and finetuned models is most likely due to the bespoke model’s overfitting.
@@ II.6 Surface segregation in high-entropy alloys
High-entropy alloys – containing 5 or more metals in roughly equimolar composition – often possess excellent mechanical properties [51], and perform well as heterogeneous catalysts [52]. Investigating their properties and exploring efficiently their composition space requires versatile models that can handle a high degree of chemical diversity. We replicate some of the simulations in Ref. 50, using PET-MAD to study the segregation of different elements at the (111) surface of the CoCrFeMnNi alloy, a prototypical HEA [51]. Modeling of the differential surface propensity of the elements in the alloy — which is central to the applications of HEAs as heterogeneous catalysts — requires a combination of replica-exchange molecular dynamics with Monte-Carlo atom swaps (REMD/MC) to overcome the slow diffusivity, and the existence of free-energy barriers to segregation starting from a random alloy. This example allows us to demonstrate the capabilities of PET-MAD in complex computational workflows, to assess the ability of the model to describe surface effects that are often overlooked in other universal models [53], and to test finetuning over a dataset with a greater degree of chemical diversity than the other examples we consider. We use the Gibbs surface excess calculated at 800 K using the REMD/MC protocol (see Section S11.3 for computational details) as a measure of surface segregation in a CoCrFeMnNi surface slab. We compare three models: the pre-trained PET-MAD, PET trained from scratch (PET-Bespoke) and the finetuned model (PET-MAD-LoRA), where the last two were trained on a subset of the HEA25S dataset [50], recomputed with MAD dataset DFT settings for consistency (details on the subset selection are given in the Supplementary Section S11.3). The validations MAEs of all three models on energy|forces are 25.8 meV/atom|175.1 meV/Å, 14.6 meV/atom|138.3 meV/Å, and 9.4 meV/atom|124.8 meV/Å, respectively.
@@
Results on surface excess calculation are given in Figure 6. First, we note that the segregation pattern obtained from both pre-trained and finetuned models is almost identical and quantitatively matches the results from Ref. 50, which corresponds to surface layers enriched in nickel and depleted in other elements. Both these models agree qualitatively with the results of the PET-Bespoke model, but there is a more pronounced quantitative difference in the surface excess values. In interpreting this discrepancy, one should keep in mind that, contrary to other examples, we only recomputed about 2000 structures out of the 30’000 in the HEA25S, and so a bespoke model incurs a high risk of overfitting – which is also consistent with the learning curves reported in the SI. This case study demonstrates that pre-trained PET-MAD is capable of capturing all effects relevant to surface segregation in HEAs and providing advanced sampling capabilities of the HEA chemical and structural space without being specifically trained on HEA data.
@@
Figure 7: Radial distribution functions $g$ of O-O and O-H pairs (top) and heat capacity $C_{V}$ of liquid water at 298 K and 1 atm, calculated using PET-MAD, a finetuned version of PET-MAD, and a bespoke PET model, as a function of the number of beads employed in the PIMD simulation (bottom). Results on the heat capacity obtained with the empirical q-TIP4P/F forcefield (reproduced from Ref. 54) are also included for comparison.
@@ II.7 Quantum nuclear effects in liquid water
Water is a widely studied system because of its high biological, environmental and technological significance, as well as the many anomalies in its physical properties, which can be traced to the strong hydrogen bonds that form local tetrahedral motifs. Due to its light hydrogen atoms, accounting for nuclear quantum effects (NQEs) is fundamental in order to extract accurate observables for liquid water, even at room temperature [55]. In this example, we employ path integral molecular dynamics (PIMD) to model nuclear quantum effects in the simulation of a medium-sized periodic system consisting of 128 water molecules (384 atoms).
We first evaluate the O-O and O-H pair correlation functions, which report on inter-molecular structural correlations, and show the extent of intra-molecular quantum fluctuations, respectively. Additionally, we calculate the constant-volume heat capacity at 298 K — a physical quantity that exhibits very pronounced quantum effects, and that requires the use of sophisticated path integral estimators that are notoriously difficult to converge [56]. We compare (1) the PET-MAD model, (2) a system-specific PET model trained on the dataset from Ref. 57, with energies and forces recomputed at the same level of theory used for MAD, (3) a PET-MAD model which was finetuned on the same recomputed water dataset and (4) values for a simple classical forcefield that reproduces well most of the experimental properties of water and allows us to demonstrate the slow convergence of some of the estimators [58].